# Superconductivity and Low Energy Excitations in an Attractive Hubbard Model

### v1

@@
Abstract: We study an attractive Hubbard model on bipartite lattices. In the grand canonical formalism, we prove the existence of superconducting long-range order in the ground state with the chemical potential corresponding to half filling. We also study the low-energy excitations above several ground states for the translationally invariant Hamiltonian. We prove the following: (i) The pairing excitations are gapless above the ground states when the number of fermions deviates from that of the half filling by the order of the volume. (ii) A certain class of single-fermion excitations shows a non-vanishing spectral gap above the ground state with an even number of fermions in a strong coupling and low-density regime.
@@ 1 Introduction
As is well known, Lieb proved two theorems [4] about the ground states of the Hubbard model on certain bipartite lattices. [^4] For previous results on the Hubbard model, the book [8] may be useful. Four years later, by relying on Lieb’s results, Shen and Qiu [5] proved the existence of superconducting long-range order in the ground state for certain fillings of fermions in the case of the attractive interaction on the so-called Lieb lattices. However, for attractive interactions, the corresponding Lieb’s theorem is restricted to the case of an arbitrary even number of fermions. Namely, the theorem is silent about the case of an odd number of fermions. The difference between even and odd numbers of fermions affects the statement of the theorem.
@@
As a result, we prove the existence of superconducting long-range order in the ground state of the attractive Hubbard model at half filling in the sense of the expectation value of the number of fermions. We stress that the long-range order implies the existence of a pure state which exhibits the spontaneous magnetization for the order parameter of superconductivity in the infinite-volume limit. Although this result is closely related to that by Shen and Qiu [5], we deal with the whole Fock space in the grand canonical formalism. In other words, we show that the sector with an odd number of fermions does not affect the existence of superconducting long-range order.
@@
The present paper is organized as follows: The precise definition of the attractive Hubbard model is given in Sec. 2, its Shiba transformation [7] is treated in Sec. 3, and the grand canonical formalism and the charge neutrality are explained in Sec.4. Following Kubo and Kishi [3], we prove the positivity of the superconducting correlation in Sec. 5. As shown in Sec. 6, the positivity yields some useful bounds. In order to use Lieb’s results which are restricted to the even numbers of fermions, we consider a certain fermion Fock space which is restricted to the corresponding sector in Sec. 7. In Sec. 8, we recall Lieb’s results. In Sec. 9, we prove the existence of superconducting long-range order for the attractive Hubbard model with the chemical potential corresponding to half filling on the Lieb lattices (Theorem 9.1). The long-range order implies the existence of a pure state which exhibits the spontaneous magnetization for the order parameter of superconductivity (Theorem 9.2). In Sec. 10, we prove the existence of a non-vanishing spectral gap above the ground state with an even number of fermions in a strong coupling and low-density regime for fermions for the lattice translationally invariant Hamiltonian of the attractive Hubbard model on the hypercubic lattices $\mathbb{Z}^{d}$ (Theorem 10.1). The case of Lieb lattices is treated in Sec. 11 (Theorem 11.1). In Sec. 12, we prove that the pairing excitations are gapless above the ground states when the number of fermions deviates from that of the half filling by the order of the volume for the lattice translationally invariant Hamiltonian (Theorems 12.1 and 12.2).
2 Model
@@
$\eta_{x}:=\begin{cases}1&\mbox{if \ }x\in\Lambda_{\rm A};\\ -1&\mbox{if \ }x\in\Lambda_{\rm B}\end{cases}$ (2.2)
@@
3 Shiba Transformation and Interaction Hamiltonian
@@
$H_{\rm int}^{(\Lambda)}=-\sum_{x\in\Lambda}U_{x}\Bigl(n_{x,\uparrow}-\frac{1}{2}\Bigr)\Bigl(n_{x,\downarrow}-\frac{1}{2}\Bigr).$ (3.2)
@@
4 Grand Canonical Formalism and Charge Neutrality
@@
where $\beta$ is the inverse temperature, and $Z_{\beta,\Lambda}:={\rm Tr}\exp[-\beta H^{(\Lambda)}]$. Here, we have chosen the chemical potential to be zero so that the half filling is realized for the particle density. To see this, we write
@@
5 Positivity of Superconducting Correlation
By using the spin reflection positivity [3], we show below the positivity (5.14) of the correlation of superconductivity. To begin with, note that
@@ 5 Positivity of Superconducting Correlation
where we have used the assumption $U_{x}>0$. By applying this identity to the right-hand side of (LABEL:Lieprod), we have
@@
for any $x,y\in\Lambda$. Since the fermion operators $\{c_{x,\sigma}\}$ have a real representation, we have
${\rm Tr}\;c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\cdot c_{x,\downarrow}^{\dagger}c_{y,\downarrow}\exp[-\beta H^{(\Lambda)}]\geq 0$ (5.13)
from the expression $\exp[-\beta H^{(\Lambda)}]$ of (5.9) and the antilinear map from $c_{x,\uparrow}$ to $c_{x,\downarrow}$. Therefore,
@@
6 Bounds for Superconducting Correlation
@@ 7 Restriction of Fermion Fock Space
8 Ground States for the Even Number of Fermions
@@ 9 Long Range Order of Superconductivity
$\langle[O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super}\rangle_{\beta,\Lambda}^{\rm even}\geq\frac{1}{|\Lambda|^{2}}\langle\!\langle[(S^{(1)})^{2}+(S^{(2)})^{2}]\rangle\!\rangle_{\beta,\Lambda}^{\rm even}+\frac{1}{|\Lambda|^{2}}\langle\!\langle S^{(3)}\rangle\!\rangle_{\beta,\Lambda}^{\rm even},$ (9.6)
@@
$\omega_{0,\Lambda}^{\rm even}([O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super})\geq\frac{1}{|\Lambda|^{2}}\check{\omega}_{0,\Lambda}^{\rm even}([(S^{(1)})^{2}+(S^{(2)})^{2}])+\frac{1}{|\Lambda|^{2}}\check{\omega}_{0,\Lambda}^{\rm even}(S^{(3)}),$ (9.7)
@@
As seen in the previous Sec. 8, the ground state of the repulsive Hubbard model is $2S_{\rm tot}+1$ fold degenerate. Therefore, one has
$\omega_{0,\Lambda}^{\rm even}([O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super})\geq\frac{1}{|\Lambda|^{2}}\frac{2}{3}S_{\rm tot}(S_{\rm tot}+1)$ (9.9)
@@
for the superconducting long-range order. Therefore, when $S_{\rm tot}=||\Lambda_{\rm A}|-|\Lambda_{\rm B}||/2$ is of order of the volume $|\Lambda|$, the ground state $\omega_{0,\Lambda}(\cdots)$ shows the long-range order at the half filling.
This theorem yields the existence of the spontaneous magnetization [2] for the superconducting order parameter $O_{\rm super}^{(\Lambda)}$ in the infinite-volume limit. In order to prove this statement, we write $H^{(\Lambda)}(B)$ for the Hamiltonian with a symmetry breaking field defined by
@@
for any observable $\mathcal{A}$. Here, $\{\Phi_{j}^{(\Lambda)}(B)\}_{j=1}^{m}$ is an orthonormal basis of the $m$-fold degenerate ground states of $H^{(\Lambda)}(B)$. We define the unitary operator $U_{\theta}^{(\Lambda)}:=\exp(i\theta n^{(\Lambda)})$ for $\theta\in[0,2\pi)$. Since the Hamiltonian $H^{(\Lambda)}$ without the external field commutes with the total number $n^{(\Lambda)}$ of fermions, we can choose an orthonormal basis of the ground states of $H^{(\Lambda)}$ so that the basis $\{\Phi^{(\Lambda)}_{j}\}_{j=1}^{\ell}$ consists of eigenstates of $n^{(\Lambda)}$, and thus of $U_{\theta}^{(\Lambda)}$ for all $\theta$. Following Theorem 7.1 in Ref. [2], we prove the existence of a spontaneous magnetization.
@@
Under Assumption 9.1, there appears a sponteneous magnetization of superconductivity in the following sense:
@@
We stress that this result implies the existence of a pure state which exhibits the spontaneous magnetization for the order parameter of superconductivity in the infinite-volume limit.
Proof of Theorem 9.2: From the superconducting long-range order (9.10), we may choose a ground state that exhibits the long-range order, namely, for each $\Lambda$ there exists $i$ such that
@@
$\Psi_{i}^{(\Lambda)}:=\frac{1}{\sqrt{2}}\Bigl(\Phi_{i}^{(\Lambda)}+\frac{O_{\Lambda}\Phi_{i}^{(\Lambda)}}{\bigl\|O_{\Lambda}\Phi_{i}^{(\Lambda)}\bigr\|}\Bigr).$
@@ 10 Spectral Gap of Single-Fermion Excitations
In this section, we require lattice translational invariance for the Hamiltonian $H^{(\Lambda)}$. More precisely, we first consider the usual hypercubic lattice $\mathbb{Z}^{d}$ of dimension $d=1,2,3,\ldots$, and then we will discuss the case of Lieb lattice. Therefore, for all the non-vanishing hopping amplitudes, we choose $t_{x,y}=t$ with a positive constant $t$, and for the coupling constants of the interactions, $U_{x}=U$ at all the lattice sites $x\in\Lambda\subset\mathbb{Z}^{d}$ with a positive constant $U$. Then, the translationally invariant Hamiltonian $H^{(\Lambda)}$ is written
$H^{(\Lambda)}=t\sum_{\sigma\in\{\uparrow,\downarrow\}}\sum_{x\in\Lambda}\sum_{y\in\Lambda:|x-y|=1}c_{x,\sigma}^{\dagger}c_{y,\sigma}-U\sum_{x\in\Lambda}\left(n_{x,\uparrow}-\frac{1}{2}\right)\left(n_{x,\downarrow}-\frac{1}{2}\right)+\mu\sum_{\sigma\in\{\uparrow,\downarrow\}}\sum_{x\in\Lambda}n_{x,\sigma}.$ (10.1)
@@
Therefore, we have $\mathcal{A}_{\uparrow}\mathcal{A}_{\uparrow}^{\dagger}+\mathcal{A}_{\uparrow}^{\dagger}\mathcal{A}_{\uparrow}=1$. From this relation, we assume
$\langle\Phi_{\rm GS},\mathcal{A}_{\uparrow}^{\dagger}\mathcal{A}_{\uparrow}\Phi_{\rm GS}\rangle\geq\frac{1}{2}\|\Phi_{\rm GS}\|^{2}$ (10.5)
for the ground state $\Phi_{\rm GS}$. Clearly, this assumption does not hold when all the lattice sites are already occupied by fermions with up spin in a ground state $\Phi_{\rm GS}$, if it exists. But we will consider the regime of low density of fermions with up spin. Actually, we can expect $\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|\approx 1$ in the regime, while $\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|\approx 0$ is expected in the high density regime.
@@
Thus the class of the single-fermion excitations that satisfies the condition (10.5) is well defined for the low-density regime of fermions. In addition, in the infinite-volume formalism, the excitations above a ground state are constructed from their finite supports [1].
In order to prove the existence of a spectral gap for single-fermionic excitations above the ground state $\Phi_{\rm GS}$, we consider
@@
Under the assumption (10.5), the following bound is valid:
$\frac{\|(H^{(\Lambda)}-E_{\rm GS})\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}{\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}\geq\Bigl(\frac{U}{2}+\mu\Bigr)-{\sqrt{2}}{|t|}\frac{\|\tilde{A}_{\uparrow}\Phi_{\rm GS}\|}{\|\Phi_{\rm GS}\|}-\sqrt{2}U\frac{\|\mathcal{A}_{\uparrow,\downarrow}\Phi_{\rm GS}\|}{\|\Phi_{\rm GS}\|},$ (10.15)
@@
$\tilde{\mathcal{A}}_{\uparrow}:=\sum_{x\in\Lambda}\tilde{\alpha}_{x}c_{x,\uparrow}^{\dagger}\quad\mbox{with}\quad\tilde{\alpha}_{y}:=\sum_{x\in\Lambda:|x-y|=1}\alpha_{x},$ (10.16)
@@
Proof: The commutator on the left-hand side of (10.14) is computed as follows:
@@
$=$ $t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda}\sum_{z\in\Lambda:|z-y|=1}[c_{y,\uparrow}^{\dagger}c_{z,\uparrow},c_{x,\uparrow}^{\dagger}]-U\sum_{x\in\Lambda}\alpha_{x}[(n_{x,\uparrow}-1/2),c_{x,\uparrow}^{\dagger}](n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda:|x-y|=1}[c_{y,\uparrow}^{\dagger}c_{x,\uparrow},c_{x,\uparrow}^{\dagger}]-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda:|x-y|=1}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $t\sum_{y\in\Lambda}\tilde{\alpha}_{y}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $\Bigl(\frac{U}{2}+\mu\Bigr)\mathcal{A}_{\uparrow}+t\sum_{y\in\Lambda}\tilde{\alpha}_{y}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}n_{x,\downarrow},$

### v2

@@
Abstract: We study an attractive Hubbard model on bipartite lattices. In the grand canonical formalism, we prove the existence of superconducting long-range order in the ground state on Lieb lattices with the chemical potential corresponding to half filling. We also study the low-energy excitations above several ground states for the translationally invariant Hamiltonian. We prove the following: (i) The pairing excitations are gapless above the ground states when the number of fermions deviates from that of the half filling by the order of the volume. (ii) A certain class of single-fermion excitations shows a non-vanishing spectral gap above the ground state with an even number of fermions in a strong coupling and low-density regime.
Keywords: Hubbard model; superconductivity; low energy excitation; long-range order; spin reflection positivity; pure state
@@ 1 Introduction
As is well known, Lieb proved two theorems [7] about the ground states of the Hubbard model on certain bipartite lattices. [^4] For previous results on the Hubbard model, see [15, 9, 8]. Four years later, by relying on Lieb’s results, Shen and Qiu [11] proved the existence of superconducting long-range order in the ground state for certain fillings of fermions in the case of the attractive interaction on the so-called Lieb lattices. To the best of our knowledge, their result is the only known proof of superconductivity for the spin-$1/2$ Hubbard model. However, for attractive interactions, the corresponding Lieb’s theorem is restricted to the case of an arbitrary even number of fermions. Namely, the theorem is silent about the case of an odd number of fermions. The difference between even and odd numbers of fermions affects the statement of the theorem.
@@
As a result, we prove the existence of superconducting long-range order in the ground state of the attractive Hubbard model on Lieb lattices. Here the system is at half filling in the sense of the expectation value of the number of fermions. We stress that the long-range order implies the existence of a pure state which exhibits the non-vanishing order parameter of superconductivity in the infinite-volume limit. Although this result is closely related to that by Shen and Qiu [11], we deal with the whole Fock space in the grand canonical formalism. In other words, we show that the sector with an odd number of fermions does not affect the existence of superconducting long-range order.
@@
The present paper is organized as follows: The precise definition of the attractive Hubbard model is given in Sec. 2. The Shiba transformation [13], the grand canonical formalism and the charge neutrality are also explained in this section. Following Kubo and Kishi [6], we prove the positivity of the superconducting correlation in Sec. 3.1. As shown in Sec. 3.2, the positivity yields some useful bounds. In order to use Lieb’s results which are restricted to the even numbers of fermions, we consider a certain fermion Fock space which is restricted to the corresponding sector in Sec. 4. In Sec. 5, we prove the existence of superconducting long-range order for the attractive Hubbard model with the chemical potential corresponding to half filling on the Lieb lattices (Theorem 5.1). The long-range order implies the existence of a pure state which exhibits the non-vanishing superconducting order parameter (Theorem 5.3). In Sec. 6, we prove the existence of a non-vanishing spectral gap above the ground state with an even number of fermions in a strong coupling and low-density regime for fermions for the lattice translationally invariant Hamiltonian of the attractive Hubbard model on the hypercubic lattices $\mathbb{Z}^{d}$ (Theorem 6.1). The case of Lieb lattices is also treated (Theorem 6.4). In Sec. 7, we prove that the pairing excitations are gapless above the ground states when the number of fermions deviates from that of the half filling by the order of the volume for the lattice translationally invariant Hamiltonian (Theorems 7.1 and 7.2).
2 Model and Properties
@@
$\eta_{x}:=\begin{cases}1&\mbox{if \ }x\in\Lambda_{\rm A},\\ -1&\mbox{if \ }x\in\Lambda_{\rm B}\end{cases}$ (2.2)
@@
2.1 Shiba Transformation and Interaction Hamiltonian
@@
$H_{\rm int}^{(\Lambda)}=-\sum_{x\in\Lambda}U_{x}\Bigl(n_{x,\uparrow}-\frac{1}{2}\Bigr)\Bigl(n_{x,\downarrow}-\frac{1}{2}\Bigr)$ (2.4)
@@
2.2 Grand Canonical Formalism and Charge Neutrality
@@
where the trace is taken over the fermion Fock space, $\beta$ is the inverse temperature, and $Z_{\beta,\Lambda}:={\rm Tr}\exp[-\beta H^{(\Lambda)}]$. Here, we have chosen the chemical potential to be zero so that the half filling is realized for the particle density. To see this, we write
@@
3 Positivity and Bounds for the Correlation
In this section, we derive a useful bound on the superconducting correlation function. For this purpose, we first establish positivity of the correlation in the next subsection.
3.1 Positivity of Superconducting Correlation
By using the spin reflection positivity [6], we show below the positivity (3.14) of the correlation of superconductivity. The result follows directly from the arguments in [6], but we provide a proof for the readers’ convenience. To begin with, note that
@@ 3.1 Positivity of Superconducting Correlation
where we have used the assumption $U_{x}>0$. By applying this identity to the right-hand side of (3.4), we have
@@
for any $x,y\in\Lambda$. We introduce an isomorphism $\theta$ from the algebra on the up-spin Fock space to that on the down-spin Fock space. In particular, $\theta(c_{x,\uparrow}^{\dagger})=c_{x,\downarrow}^{\dagger}$ and $\theta(c_{x,\uparrow})=c_{x,\downarrow}$. Then we deduce from the expression $\exp[-\beta H^{(\Lambda)}]$ of (3.9) that
${\rm Tr}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\cdot c_{x,\downarrow}^{\dagger}c_{y,\downarrow}\exp(-\beta H^{(\Lambda)})\right]=e^{\beta\mathcal{C}_{0}}\lim_{M\nearrow\infty}\int d\tilde{\mu}\,\mathrm{Tr}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\theta(c_{x,\uparrow}^{\dagger}c_{y,\uparrow})\prod_{i=1}^{M}\alpha^{(M)}_{i,\uparrow}\theta(\overline{\alpha^{(M)}_{i,\uparrow}})\right$],
where $d\tilde{\mu}:=\prod_{x\in\Lambda}\prod_{\ell=1}^{M}d\mu(k_{x,\ell})$. Since the fermion operators $\{c_{x,\sigma}\}$ have a real representation and all the $\alpha^{(M)}_{i,\sigma}$ consist of even operators, we have
$\mathrm{Tr}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\theta(c_{x,\uparrow}^{\dagger}c_{y,\uparrow})\prod_{i=1}^{M}\alpha^{(M)}_{i,\uparrow}\theta(\overline{\alpha^{(M)}_{i,\uparrow}})\right$] $=\mathrm{Tr}_{\uparrow}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\prod_{i=1}^{M}\alpha^{(M)}_{i,\uparrow}\right]\cdot\mathrm{Tr}_{\downarrow}\left[\theta\left(c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\prod_{i=1}^{M}\overline{\alpha^{(M)}_{i,\uparrow}}\right)\right$]
$=\left|\mathrm{Tr}_{\uparrow}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\prod_{i=1}^{M}\alpha^{(M)}_{i,\uparrow}\right]\right|^{2},$
where $\mathrm{Tr}_{\sigma}$ denotes the trace over the Fock space of $\sigma$-spin fermions. This proves
${\rm Tr}\left[c_{x,\uparrow}^{\dagger}c_{y,\uparrow}\cdot c_{x,\downarrow}^{\dagger}c_{y,\downarrow}\exp(-\beta H^{(\Lambda)})\right]\geq 0,$ (3.13)
which means that
@@
3.2 Bounds for Superconducting Correlation
@@ 4 Restriction of Fermion Fock Space
Ground States for an Even Number of Fermions
@@ 5 Long Range Order of Superconductivity
$\langle[O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super}^{(\Lambda)}\rangle_{\beta,\Lambda}^{\rm even}\geq\frac{1}{|\Lambda|^{2}}\langle\!\langle[(S^{(1)})^{2}+(S^{(2)})^{2}]\rangle\!\rangle_{\beta,\Lambda}^{\rm even}+\frac{1}{|\Lambda|^{2}}\langle\!\langle S^{(3)}\rangle\!\rangle_{\beta,\Lambda}^{\rm even},$ (5.6)
@@
$\omega_{0,\Lambda}^{\rm even}([O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super}^{(\Lambda)})\geq\frac{1}{|\Lambda|^{2}}\check{\omega}_{0,\Lambda}^{\rm even}([(S^{(1)})^{2}+(S^{(2)})^{2}])+\frac{1}{|\Lambda|^{2}}\check{\omega}_{0,\Lambda}^{\rm even}(S^{(3)}),$ (5.7)
@@
As seen in the previous section, the ground state of the repulsive Hubbard model is $2S_{\rm tot}+1$ fold degenerate. Therefore, one has
$\omega_{0,\Lambda}^{\rm even}([O_{\rm super}^{(\Lambda)}]^{\dagger}O_{\rm super}^{(\Lambda)})\geq\frac{1}{|\Lambda|^{2}}\frac{2}{3}S_{\rm tot}(S_{\rm tot}+1)$ (5.9)
@@
for the superconducting long-range order. Therefore, when $S_{\rm tot}=||\Lambda_{\rm A}|-|\Lambda_{\rm B}||/2$ is of order of the volume $|\Lambda|$, the ground state $\omega_{0,\Lambda}(\cdots)$ shows the long-range order at half filling.
This theorem implies that there exists a pure ground state with the non-vanishing superconducting order parameter $O_{\rm super}^{(\Lambda)}$ in the infinite-volume limit [5]. In order to prove this statement, we write $H^{(\Lambda)}(B)$ for the Hamiltonian with a symmetry breaking field defined by
@@
for any observable $\mathcal{A}$. Here, $\{\Phi_{j}^{(\Lambda)}(B)\}_{j=1}^{m}$ is an orthonormal basis of the $m$-fold degenerate ground states of $H^{(\Lambda)}(B)$. We define the unitary operator $U_{\theta}^{(\Lambda)}:=\exp(i\theta n^{(\Lambda)})$ for $\theta\in[0,2\pi)$. Since the Hamiltonian $H^{(\Lambda)}$ without the external field commutes with the total number $n^{(\Lambda)}$ of fermions, we can choose an orthonormal basis of the ground states of $H^{(\Lambda)}$ so that the basis $\{\Phi^{(\Lambda)}_{j}\}_{j=1}^{\ell}$ consists of eigenstates of $n^{(\Lambda)}$, and thus of $U_{\theta}^{(\Lambda)}$ for all $\theta$.
@@
Under Assumption 5.2, it holds that
@@
We stress that this result implies the existence of a pure state exhibiting spontaneous symmetry breaking associated with the superconducting order parameter in the infinite-volume limit.
Proof of Theorem 5.3: The proof follows from Theorem 7.1 in Ref. [5], which is based on [2, 3]. From the superconducting long-range order (5.10), we may choose a ground state that exhibits the long-range order, namely, for each $\Lambda$ there exists $i$ such that
@@
$\Psi_{i}^{(\Lambda)}:=\frac{1}{\sqrt{2}}\left(\Phi_{i}^{(\Lambda)}+\frac{O_{\Lambda}\Phi_{i}^{(\Lambda)}}{\bigl\|O_{\Lambda}\Phi_{i}^{(\Lambda)}\bigr\|}\right).$
@@ 6 Spectral Gap of Single-Fermion Excitations
6.1 The Case of Hypercubic Lattice
From now on, we will deal with excitations of the Hubbard model. In this section, we require lattice translational invariance for the Hamiltonian $H^{(\Lambda)}$. More precisely, we first consider the usual hypercubic lattice $\mathbb{Z}^{d}$ of dimension $d=1,2,3,\ldots$, and then we will discuss the case of Lieb lattice. Therefore, for all the non-vanishing hopping amplitudes, we choose $t_{x,y}=-t$ with a positive constant $t$, and for the coupling constants of the interactions, $U_{x}=U$ at all the lattice sites $x\in\Lambda\subset\mathbb{Z}^{d}$ with a positive constant $U$. Then, the translationally invariant Hamiltonian $H^{(\Lambda)}$ is written
$H^{(\Lambda)}=-t\sum_{\sigma\in\{\uparrow,\downarrow\}}\sum_{\begin{subarray}{c}x,y\in\Lambda\\ |x-y|=1\end{subarray}}c_{x,\sigma}^{\dagger}c_{y,\sigma}-U\sum_{x\in\Lambda}\left(n_{x,\uparrow}-\frac{1}{2}\right)\left(n_{x,\downarrow}-\frac{1}{2}\right)+\mu\sum_{\sigma\in\{\uparrow,\downarrow\}}\sum_{x\in\Lambda}n_{x,\sigma}.$ (6.1)
@@ 6.1 The Case of Hypercubic Lattice
Therefore, we have $\mathcal{A}_{\uparrow}\mathcal{A}_{\uparrow}^{\dagger}+\mathcal{A}_{\uparrow}^{\dagger}\mathcal{A}_{\uparrow}=1$. From this relation, we restrict our attention to single-fermion excitations $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ satisfying
$\langle\Phi_{\rm GS},\mathcal{A}_{\uparrow}^{\dagger}\mathcal{A}_{\uparrow}\Phi_{\rm GS}\rangle\geq\frac{1}{2}\|\Phi_{\rm GS}\|^{2}.$ (6.5)
As we will see in Example 6.2, such an excitation can be constructed in the low-density regime. Clearly, this relation does not hold when all the lattice sites are already occupied by fermions with up spin in a ground state $\Phi_{\rm GS}$, if it exists. But we will consider the regime of low density of fermions with up spin. Actually, we can expect $\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|\approx 1$ in the regime, while $\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|\approx 0$ is expected in the high density regime.
Now we define the excitation energy associated with the state $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ by
$\Delta(\mathcal{A}_{\uparrow}\Phi_{\rm GS}):=\frac{\|(H^{(\Lambda)}-E_{\rm GS})\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}{\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}.$
Our result in this subsection is as follows:
Theorem 6.1
Let $\Phi_{\rm GS}$ be the ground state of the translationally invariant Hamiltonian $H^{(\Lambda)}$ of (6.1) on the lattice $\Lambda\subset\mathbb{Z}^{d}$ with an even number $N>0$ of fermions, and let $\mathcal{A}_{\uparrow}$ be given by (6.2). Assume that the filling factor $\nu_{\downarrow}:=N_{\downarrow}/|\Lambda|$ for the number $N_{\downarrow}$ of fermions with down spin satisfies $\nu_{\downarrow}<1/8$. Then, for a sufficiently small $|t|/U$, the excitation energy $\Delta(\mathcal{A}_{\uparrow}\Phi_{\rm GS})$ is bounded from below by a positive constant if the state $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ satisfies the norm condition (6.5).
In other words, adding a single fermion costs a strictly positive energy above the ground state. In this paper, we also refer to $\Delta(\mathcal{A}_{\uparrow}\Phi_{\rm GS})$ as a spectral gap of single-fermion excitations.
As in the following Example 6.2, we may take $\mathcal{A}_{\uparrow}=c^{\dagger}_{x_{0},\uparrow}$ so that (6.5) holds in the low-density regime. Following Lieb’s theorem [7], for any even number $N>0$ of the fermions, the ground state is unique and shows spin singlet. Therefore, the number $N_{\uparrow}$ of fermions with up spin satisfies $N_{\uparrow}=N_{\downarrow}$ and $N=N_{\uparrow}+N_{\downarrow}$. Clearly, the filling condition $\nu_{\downarrow}<1/8$ can be realized.
@@
Thus the class of the single-fermion excitations that satisfies the condition (6.5) is well-defined for the low-density regime of fermions. In addition, in the infinite-volume formalism, the excitations above a ground state are constructed from their finite supports [4].
Proof of Theorem 6.1: In order to prove the existence of a spectral gap for single-fermionic excitations above the ground state $\Phi_{\rm GS}$, we consider
@@
Under the condition(6.5), the following bound is valid:
$\frac{\|(H^{(\Lambda)}-E_{\rm GS})\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}{\|\mathcal{A}_{\uparrow}\Phi_{\rm GS}\|}\geq\Bigl(\frac{U}{2}+\mu\Bigr)-{\sqrt{2}}{|t|}\frac{\|\tilde{\mathcal{A}}_{\uparrow}\Phi_{\rm GS}\|}{\|\Phi_{\rm GS}\|}-\sqrt{2}U\frac{\|\mathcal{A}_{\uparrow,\downarrow}\Phi_{\rm GS}\|}{\|\Phi_{\rm GS}\|},$ (6.15)
@@
$\tilde{\mathcal{A}}_{\uparrow}:=\sum_{x\in\Lambda}\tilde{\alpha}_{x}c_{x,\uparrow}^{\dagger}\quad\mbox{with}\quad\tilde{\alpha}_{x}:=\sum_{y\in\Lambda:|x-y|=1}\alpha_{y},$ (6.16)
@@
Proof of Lemma: The commutator on the left-hand side of (6.14) is computed as follows:
@@
$=$ $-t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda}\sum_{z\in\Lambda:|z-y|=1}[c_{y,\uparrow}^{\dagger}c_{z,\uparrow},c_{x,\uparrow}^{\dagger}]-U\sum_{x\in\Lambda}\alpha_{x}[(n_{x,\uparrow}-1/2),c_{x,\uparrow}^{\dagger}](n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $-t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda:|x-y|=1}[c_{y,\uparrow}^{\dagger}c_{x,\uparrow},c_{x,\uparrow}^{\dagger}]-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $-t\sum_{x\in\Lambda}\alpha_{x}\sum_{y\in\Lambda:|x-y|=1}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $-t\sum_{y\in\Lambda}\tilde{\alpha}_{y}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}(n_{x,\downarrow}-1/2)+\mu\mathcal{A}_{\uparrow}$
$=$ $\Bigl(\frac{U}{2}+\mu\Bigr)\mathcal{A}_{\uparrow}-t\sum_{y\in\Lambda}\tilde{\alpha}_{y}c_{y,\uparrow}^{\dagger}-U\sum_{x\in\Lambda}\alpha_{x}c_{x,\uparrow}^{\dagger}n_{x,\downarrow},$